OpenStudy (anonymous):

of the 8760 hrs ina yr, one tv was on for 2190 hrs. what percent is this

OpenStudy (anonymous):

2190/8760 * 100 = 25%

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

2190/8760 = 1/4 = 25%